# Severity Prediction in Mental Health: LLM-based Creation, Analysis, Evaluation of a Novel Multilingual Dataset

### v1

@@ 4.5 Resources and cost of experiments
The total cost for running the GPT3.5 experiments via its API was less than \$30, showing the minimal resources required to conduct them, without the need for expensive GPU infrastructure or fine-tuning. For example, we experimented with MentaLLaMA-chat-7B Yang et al. (2024), which is a 27GB model. Its size makes it prohibitive when limited resources are available. Eventually we ended up using the quantized version, which did not output meaningful results. Our methodology for utilizing resources is potentially promising for extending medical data sets in English into other languages as well.
@@ 5 Discussion
Despite the hype of using LLMs in mental healthcare for commercial applications, there is an increasing need for more research to understand the pitfalls and dangers of using LLMs across languages, diseases, and patient populations for mental health care. Considering the variation in the performance of the LLM across languages, there is a need for utmost precautions not to rely on LLMs solely in any healthcare setting. As stated by Stade et al. (2024), mental health diagnosis should never be left to automatic systems only and should never replace human professionals due to possible errors and/or misdiagnosis illustrated above.
LLMs exhibit considerable variability in performance across languages, despite being evaluated on the same translated dataset. This inconsistency underscores the complexities inherent in multilingual mental health support, where language-specific nuances and mental health data coverage can effect the accuracy of the models. We observe performance gains in both tasks with stronger models and by translating to a specific target language (e.g., in German, for depression severity) before detection. On the other hand, we also observe that translating to Turkish yields the worst performance in both tasks, which could be explained by the limited number of Turkish resources in health, especially in the domain of mental health Çöltekin et al. (2023).
@@ 6 Conclusion
We presented a novel multilingual dataset designed to evaluate the ability of LLMs to predict the severity of mental health conditions in user-generated posts. Using English as the source language, we translated the posts to French, German, Turkish, Greek, Finnish, and Portuguese, using LLM-based machine translation. Then, we prompted the LLM to predict the severity of two mental health conditions across all languages. Our findings show that the LLM performance varies depending on the language. Also, the best performance is not necessarily achieved in the source (English) language. Although not tested explicitly in low resource languages, our method can potentially be useful for such languages, due to its cost-effective nature. Next steps comprise the application of our methodology to a broader range of languages (including a variety of low resource languages) and adding datasets for more mental health tasks and from different social media platforms.
@@
Creating a multilingual dataset requires significant resources. In this work, we used a popular LLM like GPT3.5 to translate posts. Translating using only an LLM and not having an expert or native-language human resources introduces a small loss of information that in some cases affects the final results. Additionally, cultural differences are not “transferred" correctly when using machine translation.
@@
Automatically assessing the performance of LLMs is inherently challenging. To evaluate performance, we search for the assigned label within the LLM’s output. If no label is detected, we default to assigning the minimum label—class 0 for the depression dataset and class 0 (not suicidal) for the suicide dataset.

### v2

@@ 4.5 Resources and cost of experiments
The total cost for running the GPT3.5 experiments via its API was less than \$30, showing the minimal resources required to conduct them, without the need for expensive GPU infrastructure or fine-tuning. For example, we experimented with MentaLLaMA-chat-7B Yang et al. (2024), which is a 27GB model. Its size makes it prohibitive when limited resources are available. Eventually, we ended up using the quantized version, which did not output meaningful results. Our methodology for utilizing resources is potentially promising for extending the medical data sets (mostly in English) into other languages as well.
@@ 5 Discussion
Despite the hype of using LLMs in mental healthcare for commercial applications, there is an increasing need for more research to understand the pitfalls and dangers of using LLMs across languages, diseases, and patient populations for mental health care. Considering the variation in the performance of the LLMs across languages, there is a need for utmost precautions not to rely on them solely in any healthcare setting. As stated by Stade et al. (2024), mental health diagnosis should never be left to automatic systems; human professionals are critical due to the possibility of errors and misdiagnoses, as illustrated above.
LLMs exhibit considerable variability in performance across languages, despite being evaluated on the same translated dataset. This inconsistency underscores the complexities inherent in multilingual mental health support, where language-specific nuances and mental health data coverage can effect the accuracy of the models. We observe performance gains in both tasks with stronger models and by translating to a specific target language (e.g., in German, for depression severity) before detection. On the other hand, translating to Turkish yields the worst performance in both tasks, which could be explained by the limited number of Turkish resources in health, especially in the domain of mental health Çöltekin et al. (2023).
@@ 6 Conclusion
We presented a novel multilingual dataset designed to evaluate the ability of LLMs to predict the severity of mental health conditions in user-generated posts. Using English as the source language, we translated the posts to French, German, Turkish, Greek, Finnish, and Portuguese, using LLM-based machine translation. Then, we prompted the LLM to predict the severity of two mental health conditions across all languages. Our results indicate that the LLM performance varies depending on the language. Also, the best performance is not necessarily achieved in the source (English) language. Tested on low-resource languages (i.e., Greek and Turkish on medical domain), our method is useful for other languages due to its cost-effective nature. Next steps comprise the application of our methodology to a broader range of languages (including a variety of low-resource languages) and adding datasets for more mental health tasks.
@@
Creating a multilingual dataset requires significant resources. In this work, we used a popular LLM like GPT3.5 to translate posts. Translating using only an LLM and not having an expert or native-language human resources introduces a small loss of information that in some cases affects the final results. Additionally, cultural differences are not “transferred” correctly when using machine translation.
@@
Automatically assessing the performance of LLMs is inherently challenging. To evaluate performance, we search for the assigned label within the LLM’s output. If no label is detected, we assign the minimum label—class 0 for the depression dataset and class 2 for the suicide dataset (Indicator).
@@
Appendix A Appendix
Next, we present some examples of misprediction when using the original post, translated and back-translated text. All examples come from the C-ssrs dataset.
A.1 Translation mispredicted, back-translation ok
Original post: “Im sure that no ones lying to you. A man I know just succeeded in committing Suicide and I can assure you that he was a ghost of himself, consumed by major Mental Depression for months. Everyone cares, and everyone wishes he hadnt done it. I hope you can find a way to carry on, and I hope that if you give it some time you can find healing and happiness in your life.”
Ground-truth: 2. Indicator
Prediction: 2. Indicator
Here the original post was correctly classified as “2. Indicator”.
Finnish translation: “Olen varma, ettei kukaan valehtele sinulle. Mies, jonka tunnen, onnistui juuri tekemään itsemurhan, ja voin vakuuttaa, että hän oli varjo entisestään, syötyään pahoista mielenterveysongelmista kuukausia. Kaikki välittävät, ja kaikki toivovat, ettei hän olisi tehnyt sitä. Toivon, että löydät tavan jatkaa eteenpäin, ja toivon, että jos annat sille aikaa, voit löytää parantumista ja onnea elämääsi.”
Prediction: 3. Ideation
Back-translation to English: “I am sure that no one is lying to you. The man I know just managed to commit suicide, and I can assure you that he was a shadow of his former self, having suffered from severe mental health issues for months. Everyone cares, and everyone wishes he hadn’t done it. I hope you find a way to move forward, and I hope that if you give it time, you can find healing and happiness in your life.”
Prediction with back-translation: 2. Indicator
In this example we observe that the back-translation helped to correct the prediction.
A.2 Translation ok, back-translation misprediction
Original post: “Dont feel guilty. Better than doing nothing, get someone(his family or a friend) to show him that he is cared for. If you get a chance you could go personally and help him. For now keep him encouraged. Try to cheer him up and distract him from these thoughts. I understand how you feel. I also get such thoughts. If you need a friend, Ill be there for you(you can PM me). Dont waste yourself. You are a blessing to this world.Your family will feel Tired sad. BTW, its better you try taking that anti-depressing medication. Please throw away the gun. True. Everyone has a guardian angel.”
Ground-truth: 1. Supportive
Prediction: 2. Indicator
Here the original post was misclassified as “2. Indicator” while it belongs to “1. Supportive”.
[…]
Πρεδιςτιον ωιτη Γρεεϰ τρανςλατιον: 1. Συππορτιε
Βαςϰ-τρανςλατιον το Ενγλιςη: ``Δο νοτ φεελ γυιλτψ. Ινςτεαδ οφ δοινγ νοτηινγ, φινδ ςομεονε (τηειρ φαμιλψ ορ α φριενδ) το ςηοω τηεμ τηατ ψου ςαρε. Ιφ ψου ηαε τηε οππορτυνιτψ, ψου ςαν γο ιν περςον ανδ ηελπ τηεμ. Φορ νοω, ϰεεπ τηεμ ενςουραγεδ. Τρψ το λιφτ τηειρ ςπιριτς ανδ διςτραςτ τηεμ φρομ τηοςε τηουγητς. Ι υνδερςτανδ ηοω ψου φεελ. Ι αλςο ηαε ςυςη τηουγητς. Ιφ ψου νεεδ α φριενδ, Ι ωιλλ βε ηερε φορ ψου (ψου ςαν μεςςαγε με). Δο νοτ ωαςτε ψουρςελφ. Ψου αρε α βλεςςινγ το τηις ωορλδ. Ψουρ φαμιλψ ωιλλ φεελ τιρεδ ανδ ςαδ. Αλςο, ιτ ις βεττερ το τρψ ταϰινγ τηις αντιδεπρεςςαντ μεδιςατιον. Πλεαςε τηροω τηε γυν αωαψ. Τρυε. Εερψ περςον ηας α γυαρδιαν ανγελ.᾽
Πρεδιςτιον ωιτη βαςϰ-τρανςλατεδ: 2. Ινδιςατορ.
Ωε οβςερε τηατ τηε τεξτ ωας ινςορρεςτλψ πρεδιςτεδ, ωηιλε τηε Γρεεϰ τρανςλατιον ωας πρεδιςτεδ ςορρεςτλψ.